# The Poisson Midpoint Method for Langevin Dynamics: Provably Efficient Discretization for Diffusion Models

### v1

@@ 1 Introduction
The task of sampling from a target distribution is central to Bayesian inference, generative modeling, differential privacy and theoretical computer science [48, 20, 17, 25]. Sampling algorithms, based on the discretization of a stochastic differential equation (SDE) called the Langevin Dynamics, are widely used. The straight forward time discretization (i.e, Euler Maruyama discretization) of Langevin dynamics, called Langevin Monte Carlo (LMC), is popular due to its simplicity. The convergence properties of LMC have been studied extensively in the literature under various conditions on the target distribution [8, 12, 11, 45, 15, 33, 7, 1, 6, 9, 16, 30, 53, 5]. LMC can suffer from slow convergence to the target distribution, and often requires a large number of steps with a very fine time discretization (i.e, small step-size), making it prohibitively expensive.
@@ 1.1 Prior Work
Euler Maruyama discretization of SDEs is known to be inefficient and many powerful numerical integration techniques have been studied extensively ( [23, 32, 3, 26]). However, higher order methods such as the Runge-Kutta method require the existence and boundedness of higher order derivatives of the drift. The Randomized Midpoint Method for LMC (RLMC) was introduced for strongly log-concave sampling [40] and was further explored in [50, 18]. RLMC is popular due to its simplicity, and ease of implementation and does not require higher order bounded derivatives of the drift function. However, the current theoretical results are restricted to the case of strongly log-concave sampling, whereas non-log-concave sampling is of immense practical interest.
@@ 1.2 Our Contributions
(1) We design the Poisson Midpoint Method which discretizes SDEs by approximating $K$-steps of Euler-Maruyama discretization with step-size $\frac{\alpha}{K}$ with just one step with step-size $\alpha$. We show a strong error bound between these two processes under general conditions in Theorem 1 (no assumption on mixing, smoothness etc). This is based on a Central Limit Theorem (CLT) based method in [10] to analyze stochastic approximations LMC.
@@ 2 Problem Setup
Consider an iterative, discrete time process $(X_{t})_{t\in\mathbb{N}\cup\{0\}}$ over $\mathbb{R}^{d}$, with step-size $\alpha>0$ given by:
@@
Here, $A_{\alpha}=\mathbf{I}$, $G_{\alpha}=\alpha\mathbf{I}$, $\Gamma_{\alpha}=\sqrt{2\alpha}\mathbf{I}$ and $b(\mathbf{x},t\alpha)=-\nabla F(\mathbf{x})$ for some $F:\mathbb{R}^{d}\to\mathbb{R}$. This is the Euler-Maruyama discretization of Overdamped Langevin Dynamics:
@@
Here $B_{\tau}$ is the standard Brownian motion in $\mathbb{R}^{d}$. Under mild conditions on $F$ and $\bar{X}_{0}$, $\mathsf{Law}(\bar{X}_{\tau})\stackrel{{\scriptstyle\tau\to\infty}}{{\to}}\pi^{*}$ where $\pi^{\star}(X)\propto\exp(-F(X))$ is the stationary distribution. $X_{t}$ in (1) approximates $\bar{X}_{\alpha t}$ [38, 35]. OLMC is the canonical algorithm for sampling and has been studied under assumptions such as log-concavity of $\pi^{\star}$ [8, 12, 11] or that $\pi^{\star}$ satisfies isoperimetric inequalities [45, 15, 33, 7, 1].
@@
This is the Euler-Maruyama discretization of Underdamped Langevin Dynamics (a.k.a. the Kinetic Langevin Dynamics) studied extensively in Physics[14]):
$d\bar{U}_{\tau}$ $=\bar{V}_{\tau}d\tau;\qquad d\bar{V}_{\tau}=-\gamma\bar{V}_{\tau}-\nabla F(\bar{U}_{\tau})+\sqrt{2}dB_{\tau}$ (3)
@@
See [43, 5, 4] for further details. The drift $\nabla\log p_{\tau}$ is learned via neural networks for discrete time instants $\tau_{0},\dots,\tau_{n-1}$ (usually $n=1000$). In practice, the iterations are written in the form  [^2] The time convention in the DDPM literature is reverse: $X_{t-1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{n-1-t}}(X_{t})+\sigma_{t}Z_{t}$.: $X_{t+1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{t}}(X_{t})+\sigma_{t}Z_{t}$ where $a_{t},b_{t},\sigma_{t}$ are chosen for best performance. Aside from the original choice in [20], many others have been proposed ([2, 42]). Since $A_{\alpha},G_{\alpha},\Gamma_{\alpha}$ in Equation (1) are time independent, we provide a variant of the Poisson Midpoint Method to suit DDPMs in Section A.1, along with a few other optimizations.
@@ 2.1 Technical Notes
We illustrate the Randomized Midpoint Method [40, 18, 50] by applying it to OLMC (to obtain RLMC) to motivate our method and explain why we expect a quadratic speed up. Overdamped Langevin Dynamics (2) satisfies:
@@ 2.2 The Poisson Midpoint Method
We approximate $K$ steps of $\mathsf{S}(A,G,\Gamma,b,\frac{\alpha}{K})$ with step-size $\frac{\alpha}{K}$ with one step of our algorithm $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ which has a step-size $\alpha$. The iterates of $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ are denoted by $(\tilde{X}_{tK})_{t\geq 0}$ or $(X_{t}^{\mathsf{P}})_{t\geq 0}$. Suppose $H_{t,i}\in\{0,1\}$ be any binary sequence and $Z_{tK+i}$ be a sequence of i.i.d. $\mathcal{N}(0,\mathbf{I})$ for $t,i\in\mathbb{N}\cup\{0\},0\leq i\leq K-1$. Given $\tilde{X}_{tK}$, we define the cheap interpolation:
@@
We then define the refined iterates for $10\leq i\leq K-$ 1 for a given $t$ as:
@@ 3 Main Results
Let $X_{t}$ be the iterates of $\mathsf{S}(A,G,\Gamma,\tfrac{\alpha}{K})$ and $X^{\mathsf{P}}_{t}$ be the iterates of $\mathsf{PS}(A,G,\Gamma,\alpha,K)$ with OPTION 1. Suppose that $X_{0}^{\mathsf{P}}=X_{0}$. Let $\tilde{X}_{tK+i}$ be the iterates in Equation (8). Define random variables:
@@
$F:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L$-smooth (i.e, $\nabla F$ is $L$-Lipschitz). $\mathbf{x}^{*}$ is its global minimizer.
@@
The assumptions above are very mild and standard in the literature. We do not make any assumptions regarding isoperimetry of the target distribution $\pi^{\star}(\mathbf{x})\propto\exp(-F(\mathbf{x}))$.
@@
$\quad+CL^{4}\alpha^{4}Kd^{2}T+\text{Lower Order Terms}$ (11)
The lower order terms are explicated in Equation (40). The next theorem gives a similar guarantee for ULMC and the lower order terms are explicated in Equation 60.
@@
$\quad+\tfrac{K\alpha^{7}L^{4}T^{2}}{\gamma}(d+\log K)^{2}+\text{ Lower Order Terms}$ (12)
@@
Table 1: Comparison of LMC and PLMCguarantees. LMC steps is the upper bound for $T$ for LMC to achieve the error guarantee in the referenced work. PLMC steps is the corresponding upper bound for PLMC. PLMC obtains a quadratic improvement in $\epsilon$, and improved dependence on $\frac{L}{\lambda},d$
Algorithm Reference Conditions LMC steps PLMC steps (ours)
@@ 4 Proof Sketch
$\mathbb{E}F(\tilde{X}_{(t+1)K})-\mathbb{E}F(\tilde{X}_{tK})\lesssim-\alpha\|\nabla F(\tilde{X}_{tK})\|^{2}+\alpha d+\text{ Lower order terms}$
$\implies\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim\frac{F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x})}{\alpha}+LTd+\text{ Lower order terms}$
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\lesssim L^{p-1}\mathbb{E}\frac{(F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x}))^{p}}{\alpha}+TL^{p}d^{p}(1+(\alpha LT)^{p-1})+\text{lower order terms}$
@@ 5 Experiments
We now present experiments to evaluate Poisson Midpoint Method as a training-free scheduler for diffusion models. We consider the Latent Diffusion Model (LDM) [39] for CelebAHQ 256, LSUN Churches, LSUN Bedrooms and FFHQ datasets using the official (PyTorch) codebase and checkpoints. We compare the sample quality of the Poisson Midpoint Method against established methods such as DDPM, DDIM and DPMSolver, varying the number of neural network calls (corresponding to the drift $b(\mathbf{x},t)$) used to generate a single image.
@@ 6 Conclusion
We introduce the Poisson Midpoint Method, which efficiently discretizes Langevin Dynamics and theoretically demonstrates quadratic speed up over Euler-Maruyama discretization under general conditions. We apply our method to diffusion models for image generation, and show that our method maintains the quality of 1000 step DDPM with just 50-80 neural network calls. This outperforms ODE based methods such as DPM-Solver in terms of quality, with a similar amount of compute. Future work can explore variants of Poisson midpoint method with better performance when fewer than 50 neural network calls are used.
@@ 7 Societal Impact
We thank Prateek Jain for helpful comments during the course of this research work.
@@ A.1 Pseudocode
    We take the dynamics to go backward in time (i.e, compute $X_{t-1}$ from $X_{t}$)
@@ Appendix C Proof of Theorem 1
The proof relies on the Wasserstein CLT based approach introduced in [10] to compare two discrete time stochastic processes. For the sake of clarity we consider the drift function $b(x,t)$ to be $b(x)$ (i.e, time invariant). However, the proof goes through for time varying drifts as well.
@@ C.3 Controlling KL Divergence via Wasserstein Distances
$\leq\mathbb{E}\mathcal{W}_{2}^{2}\left(\mathsf{Law}(Z_{\tau,2}),\mathsf{Law}(Z_{\tau,2}+\sqrt{2}S_{\tau}|\tilde{Z}_{0:\tau-1},X_{0})\right)+\mathbb{E}[\|B_{\tau}\|^{2}|\tilde{Z}_{0:\tau-1},X_{0}]$ (17)
@@ C.4 Bounding the Error Along the Trajectory:
    $Q:=\mathsf{Law}(Y_{Kt+i,2}+\sqrt{2}S_{Kt+i}|\tilde{Z}_{0:Kt+i-1},Z_{0,Kt+i-1},X_{0})$
@@ F.1 Bounding the Moments:
    $\quad+\|\sum_{j=0}^{i-1}A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Y_{tK+j}\|_{\mathsf{any}}$

### v2

@@ 1 Introduction
The task of sampling from a target distribution is central to Bayesian inference, generative modeling, differential privacy and theoretical computer science [48, 20, 17, 25]. Sampling algorithms, based on the discretization of a stochastic differential equation (SDE) called the Langevin Dynamics, are widely used. The straightforward time discretization (i.e., Euler Maruyama discretization) of Langevin dynamics, called Langevin Monte Carlo (LMC), is popular due to its simplicity. The convergence properties of LMC have been studied extensively in the literature under various conditions on the target distribution [8, 12, 11, 45, 15, 33, 7, 1, 6, 9, 16, 30, 53, 5]. LMC can suffer from slow convergence to the target distribution, and often requires a large number of steps with a very fine time discretization (i.e., small step-size), making it prohibitively expensive.
@@ 1.1 Prior Work
Euler Maruyama discretization of SDEs is known to be inefficient and many powerful numerical integration techniques have been studied extensively ( [23, 32, 3, 26]). However, higher order methods such as the Runge-Kutta method require the existence and boundedness of higher order derivatives of the drift. The Randomized Midpoint Method for LMC (RLMC) was introduced for strongly log-concave sampling [40] and was further explored in [50, 18]. It was shown that RLMC, under certain conditions, can sample with a larger step size for fewer steps compared to Euler Maruyama and yet obtain the same accuracy. RLMC is popular due to its simplicity, and ease of implementation and does not require higher order bounded derivatives of the drift function. However, the current theoretical results are restricted to the case of strongly log-concave sampling, whereas non-log-concave sampling is of immense practical interest.
@@ 1.2 Our Contributions
(1) We design the Poisson Midpoint Method which discretizes SDEs by approximating $K$-steps of Euler-Maruyama discretization with step-size $\frac{\alpha}{K}$ by just one step with step-size $\alpha$. We show a strong error bound between these two processes under general conditions in Theorem 1 (no assumption on mixing, smoothness etc). This is based on a Central Limit Theorem (CLT) based method in [10] to analyze stochastic approximations of LMC.
@@ 2 Problem Setup
Given a random vector $X_{0}\in\mathbb{R}^{d}$, consider an iterative, discrete time process $(X_{t})_{t\in\mathbb{N}\cup\{0\}}$ over $\mathbb{R}^{d}$, with step-size $\alpha>0$ given by:
@@
Consider Overdamped Langevin Dynamics for some $F:\mathbb{R}^{d}\to\mathbb{R}$:
@@
Here $B_{\tau}$ is the standard Brownian motion in $\mathbb{R}^{d}$. Under mild conditions on $F$ and $\bar{X}_{0}$, $\mathsf{Law}(\bar{X}_{\tau})\stackrel{{\scriptstyle\tau\to\infty}}{{\to}}\pi^{*}$ where $\pi^{\star}(X)\propto\exp(-F(X))$ is the stationary distribution.
Picking $A_{\alpha}=\mathbf{I}$, $G_{\alpha}=\alpha\mathbf{I}$, $\Gamma_{\alpha}=\sqrt{2\alpha}\mathbf{I}$ and $b(\mathbf{x},t\alpha)=-\nabla F(\mathbf{x})$ in Equation (1) gives us Euler-Maruyama discretization of Overdamped Langevin Dynamics: $X_{t}$ in (1) approximates $\bar{X}_{\alpha t}$ [38, 35]. OLMC is the canonical algorithm for sampling and has been studied under assumptions such as log-concavity of $\pi^{\star}$ [8, 12, 11] or that $\pi^{\star}$ satisfies isoperimetric inequalities [45, 15, 33, 7, 1].
@@
This choice of $A_{h},\Gamma_{h},G_{h},b(,)$ in Equation (1) gives the Euler-Maruyama discretization of Underdamped Langevin Dynamics (a.k.a. the Kinetic Langevin Dynamics) studied extensively in Physics[14]):
$d\bar{U}_{\tau}$ $=\bar{V}_{\tau}d\tau;\qquad d\bar{V}_{\tau}=-\gamma\bar{V}_{\tau}-\nabla F(\bar{U}_{\tau})+\sqrt{2\gamma}dB_{\tau}$ (3)
@@
See [43, 5, 4] for further details. Here $p_{\tau}$ is the probability density of $e^{-t}X^{*}+\sqrt{1-e^{-2t}}Z$ where $X^{*}$ is drawn from the target and $Z$ is drawn from $\mathcal{N}(0,\mathbf{I})$ independently. The drift $\nabla\log p_{\tau}$ is learned via neural networks for discrete time instants $\tau_{0},\dots,\tau_{n-1}$ (usually $n=1000$). In practice, the iterations are written in the form  [^2] The time convention in the DDPM literature is reverse: $X_{t-1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{n-1-t}}(X_{t})+\sigma_{t}Z_{t}$.: $X_{t+1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{t}}(X_{t})+\sigma_{t}Z_{t}$ where $a_{t},b_{t},\sigma_{t}$ are chosen for best performance. Aside from the original choice in [20], many others have been proposed ([2, 42]). Since $A_{\alpha},G_{\alpha},\Gamma_{\alpha}$ in Equation (1) are time independent, we provide a variant of the Poisson Midpoint Method to suit DDPMs in Section A.1, along with a few other optimizations.
@@ 2.1 Technical Notes
We first illustrate the Randomized Midpoint Method [40, 18, 50] by applying it to OLMC (to obtain RLMC) to motivate our method (the Poisson Midpoint Method) and explain why we expect a quadratic speed up shown in Section 3. Overdamped Langevin Dynamics (2) satisfies:
@@ 2.2 The Poisson Midpoint Method
We introduce the Poisson Midpoint Method (PLMC) which approximates $K$ steps of $\mathsf{S}(A,G,\Gamma,b,\frac{\alpha}{K})$ with step-size $\frac{\alpha}{K}$ with one step of which has a step-size $\alpha$. We denote this by $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ and let its iterates be denoted by $(\tilde{X}_{tK})_{t\geq 0}$ or $(X_{t}^{\mathsf{P}})_{t\geq 0}$. Suppose $H_{t,i}\in\{0,1\}$ be any binary sequence and $Z_{tK+i}$ be a sequence of i.i.d. $\mathcal{N}(0,\mathbf{I})$ for $t,i\in\mathbb{N}\cup\{0\},0\leq i\leq K-1$. Given $\tilde{X}_{tK}$, we define the interpolation:
@@
Note that this interpolation is cheap since every one of $\hat{\tilde{X}}_{tK+i}$ can be computed with just one evaluation of the function $b()$. We then define the refined iterates for $10\leq i\leq K-$ 1 for a given $t$ as:
@@
That is, the algorithm first generates the random mid-points $H_{t,i}$, computes the interpolation $\hat{\tilde{X}}_{tK+i}$ only when $H_{t,i}=1$ and then computes $b(\hat{\tilde{X}}_{tK+i},\frac{(tK+i)\alpha}{K})$ for these points. These computations are then combined to compute $\tilde{X}_{(t+1)K}$. In most applications, it is computationally easy to generate Gaussian random vectors and perform vector operations such as summation. However, the evaluation of the drift function $b()$ is expensive. Therefore, in this work, we consider the number of evaluations of the drift function as the measure of computational complexity. The following proposition establishes that each iteration of PLMC requires 2 evaluations of $b()$ in expectation.
@@ 3 Main Results
Let $X_{t}$ be the iterates of $\mathsf{S}(A,G,\Gamma,b,\tfrac{\alpha}{K})$ and $X^{\mathsf{P}}_{t}$ be the iterates of $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ with OPTION 1. Suppose that $X_{0}^{\mathsf{P}}=X_{0}$. Let $\tilde{X}_{tK+i}$ be the iterates in Equation (8). Define random variables:
@@
$F:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L$-smooth (i.e., $\nabla F$ is $L$-Lipschitz). $\mathbf{x}^{*}$ is its global minimizer.
@@
The assumptions above are very mild and standard in the literature. Specifically, Assumption 2 shows that the initialization is close to global optimum by $O(\sqrt{d})$ up to 14th moments. For instance, this is satisfied when the initialization is a standard Gaussian variable with mean $\mu$ satisfying $\|\mu-\mathbf{x}^{*}\|=O(\sqrt{d})$. Specifically this is true when $\mu=0$ and $\|\mathbf{x}^{*}\|=O(\sqrt{d})$. This is a weak assumption which is implied from common initialization assumptions in the literature as listed below. It can be replaced with the assumptions in [53, Appendix D and Lemma 27] which considers Gaussian initializations with the right variance and mean. The original randomized midpoint method work [40] considers initializing at $\mathbf{x}^{*}$ whereas [45] considers a Gaussian initialization with the right variance at a local minimum of $F$.
We do not make any assumptions regarding isoperimetry of the target distribution $\pi^{\star}(\mathbf{x})\propto\exp(-F(\mathbf{x}))$.
@@
$\quad+O(CL^{4}\alpha^{4}Kd^{2}T)$ (11)
Remark 2.
There are lower order terms hidden in the $O()$ notation. These are explicated in Equation (40) in the appendix. The next theorem gives a similar guarantee for ULMC and the lower order terms are explicated in Equation 60 in the appendix.
@@
$\quad+O(\tfrac{K\alpha^{7}L^{4}T^{2}}{\gamma}(d+\log K)^{2})$ (12)
@@
Table 1: Comparison of LMC and PLMC guarantees. LMC complexity is the upper bound on the number of drift ($b()$) evaluations to achieve the error guarantee in the referenced work. PLMC complexity is the corresponding upper bound for PLMC. PLMC obtains a quadratic improvement in $\epsilon$, and improved dependence on $\frac{L}{\lambda},d$. The bounds hold up to poly-log factors.
Algorithm Reference Conditions LMC complexity PLMC complexity
@@ 4 Proof Sketch
$\mathbb{E}F(\tilde{X}_{(t+1)K})-\mathbb{E}F(\tilde{X}_{tK})\lesssim-\alpha\|\nabla F(\tilde{X}_{tK})\|^{2}+\alpha d+o(\alpha d)$
$\implies\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim\frac{F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x})}{\alpha}+LTd+o(LTd)$
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\lesssim L^{p-1}\mathbb{E}\frac{(F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x}))^{p}}{\alpha}+TL^{p}d^{p}(1+(\alpha LT)^{p-1})$
@@ 5 Experiments
We now present experiments to evaluate Poisson Midpoint Method as a training-free scheduler for diffusion models. We consider the Latent Diffusion Model (LDM) [39] for CelebAHQ 256, LSUN Churches, LSUN Bedrooms and FFHQ datasets using the official (PyTorch) codebase and checkpoints. We compare the sample quality of the Poisson Midpoint Method against established methods such as DDPM, DDIM and DPM-Solver, varying the number of neural network calls (corresponding to the drift $b(\mathbf{x},t)$) used to generate a single image.
@@ 6 Conclusion
We introduce the Poisson Midpoint Method, which efficiently discretizes Langevin Dynamics and theoretically demonstrates quadratic speed up over Euler-Maruyama discretization under general conditions. We apply our method to diffusion models for image generation, and show that our method maintains the quality of 1000 step DDPM with just 50-80 neural network calls. This outperforms ODE based methods such as DPM-Solver in terms of quality, with a similar amount of compute. Future work can explore variants of Poisson midpoint method with better performance when fewer than 50 neural network calls are used. An interesting theoretical direction would be to derive convergence bounds for algorithms such as DDPM which have a time dependent drift function. Future work can also consider convergence rates of PLMC under conditions such as the Poincare Inequality and whenever $\nabla F$ is Hölder continuous instead of Lipschitz continuous.
@@ 7 Societal Impact
We thank Prateek Jain for helpful comments during the course of this research work. This work was done when SK was a student researcher at Google Research.
@@ A.1 Pseudocode
    We take the dynamics to go backward in time (i.e., compute $X_{t-1}$ from $X_{t}$)
@@ Appendix C Proof of Theorem 1
The proof relies on the Wasserstein CLT based approach introduced in [10] to compare two discrete time stochastic processes. For the sake of clarity we consider the drift function $b(x,t)$ to be $b(x)$ (i.e., time invariant). However, the proof goes through for time varying drifts as well.
@@ C.3 Controlling KL Divergence via Wasserstein Distances
$\leq\mathcal{W}_{2}^{2}\left(\mathsf{Law}(Z_{\tau,2}),\mathsf{Law}(Z_{\tau,2}+\sqrt{2}S_{\tau}|\tilde{Z}_{0:\tau-1},X_{0})\right)+2\mathbb{E}[\|B_{\tau}\|^{2}|\tilde{Z}_{0:\tau-1},X_{0}]$ (17)
@@ C.4 Bounding the Error Along the Trajectory:
    $Q:=\mathsf{Law}(Z_{Kt+i,2}+\sqrt{2}S_{Kt+i}|\tilde{Z}_{0:Kt+i-1},Z_{0,Kt+i-1},X_{0})$
@@ F.1 Bounding the Moments:
    $\quad+\|\sum_{j=0}^{i-1}A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}\|_{\mathsf{any}}$